# Online Resource Allocation via Static Bundle Pricing

### v1

@@ 3.2 The Unconstrained Algorithm
Observe that any buyer with value strictly larger than $w$ is always included in UAlg. In contrast, buyers whose value equals $w$ may or may not belong to UAlg, depending on the arrival order chosen by the adversary. Thus we consider the contribution from buyers with tight and non-tight values separately.
We first focus on the second term in the last inequality. We consider three cases regarding the menu entries at price $w$ we post for set $S$. In the following we carefully handle the events $\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}$ and $\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}$ to decouple the randomness in the realization and the menu.
Case 1: We offer one copy with probability $x_{S,w}/q_{S,w}$.
Thus, if $\in\textnormal{{UAlg}}_{M,I}, bb$ got the single copy at crucial value and we get (recall that $\ell_{A}(e)$ denotes the load of item $e$ due to buyers in $A$):
$\mathcal{E}_{b}\subseteq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b^{\prime}\in\mathcal{A},\ b^{\prime}\neq b}\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}\geq c(e)\right\}\eqqcolon\mathcal{F}_{b}$
Note that (from the proof of Lemma 3.4) $\mathbb{P}\!\left[\mathcal{F}_{b}\right]\leq 0.1$ since for each $e\in S,\ M,\ I$
$\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b^{\prime}\in\mathcal{A},\ b^{\prime}\neq b}\mathbbm{1}{\left\{v_{b^{\prime}}>w\right\}}\leq Z_{M,I}(e).$
Then, the second term is upper bounded by
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}w\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{F}_{b}\right\}}\right]$
@@
[…]
Case 2: We offer one copy with probability $x_{S,w}$.
Define
$\mathcal{F}_{A}\coloneqq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\geq c(e)\right\}$
Note that $\mathbb{P}\!\left[\mathcal{F}_{A}\right]\leq 0.1$ similarly to Case 1. Now it holds that for every $M$ and for every II
[…]
Case 3: We offer $\lfloor x_{S,w}\rfloor$ copies.
Define
$\mathcal{F}_{\mathcal{A}}\coloneqq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I}\setminus\mathcal{A}}(e)+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}+\left\lfloor x_{S,w}\right\rfloor\geq c(e)\right\}$
Note that $\mathbb{P}\!\left[\mathcal{F}_{\mathcal{A}}\right]\leq 0.1$ similarly to Case 1. Now it holds that for every $M$ and for every II
$\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}=w\right\}}\mathbbm{1}{\left\{b\in\textnormal{{UAlg}}_{M,I}\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\leq\ \left\lfloor x_{S,w}\right\rfloor\cdot\mathbbm{1}{\left\{\mathcal{F}_{\mathcal{A}}\right\}}$
Indeed, the left-hand side is upper bounded by $\left\lfloor x_{w}\right\rfloor$ and if $\mathbbm{1}{\left\{\mathcal{F}_{\mathcal{A}}\right\}}=0$ then no buyer in $\mathcal{A}$ can be blocked. Taking expectations and multiplying with $w$ yields that the second term is upper bounded by $0.1\cdot w\left\lfloor x_{S,w}\right\rfloor\leq 0.1\cdot w\cdot x_{S,w}$.
Combining all the previous cases and substituting back to ($\star\star$4) we get:
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]\leq\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{E}_{b}\right\}}\right]+0.1\cdot x_{S,w}\cdot w$
Now note that
$\mathcal{E}_{b}\subseteq\left\{\exists e\in S\,:\,\ell_{\textnormal{{UAlg}}_{M,I\setminus\{b\}}}(e)\geq c(e)\right\}\eqqcolon\mathcal{G}_{b}$
since the load due to previous buyers (with respect to $b$) in any item is smaller than the load due to UAlg ignoring $b$ from the instance $I$. Now noting that $\mathbb{P}\!\left[\mathcal{G}_{b}\right]\leq 0.1$ since $\ell_{\textnormal{{UAlg}}_{M,I\setminus\{b\}}}(e)$ is upper bounded by $Z_{M,I}(e)$ and using the fact that $\mathcal{G}_{b}$ is independent to $v_{b}$ we get
$\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}\cap\textsc{Blocked}_{M,I}}v_{b}\right]$ $\leq\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{A}}v_{b}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\mathcal{G}_{b}\right\}}\right]+0.1\cdot x_{S,w}\cdot w$
$\leq\sum_{b\in\mathcal{A}}\sum_{v>w}v\cdot q_{b,v}\cdot 0.1+0.1\cdot x_{S,w}\cdot w=0.1\textnormal{{FracOpt}}_{S}$
and summing over all possible $S$ yields the proof. ∎ As a result, ($\star$4) becomes
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]$ $\;\geq\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;0.1\textnormal{{FracOpt}}_{\gamma}\geq\frac{1}{40}\cdot\textnormal{{FracOpt}}_{\gamma}$
$\;\geq\;\frac{\textnormal{{FracOpt}}}{40e(10d)^{1/B}}$ ($\star$)
5 General Single-Minded Setting
We now turn to the most general case, where buyers are single-minded over arbitrary bundles. We proceed to prove Theorem˜1.2.
Approach.
Our algorithmic approach follows the same template as before: we solve the scaled-down LP (Step 1) with an appropriate choice of $\gamma$, construct the randomized menu (Step 2), and analyze the performance of the resulting posted-price mechanism.
In this setting, however, very large valuations can dominate the benchmark and invalidate the guarantees obtained by the LP-based menu alone. To handle this, we introduce an auxiliary mechanism that posts a single high-priced bundle; the final algorithm randomly mixes between these two mechanisms.
Analysis.
We select $\gamma\;=\;e(20m)^{1/(B+1)},$ and let $\textnormal{{FracOpt}}_{\gamma}$ be the value of the scaled LP. When a buyer arrives, they may desire a menu entry but fail to obtain it due to capacity conflicts with previously accepted buyers. As before, we define:
$\textsc{Blocked}_{M,I}=\{b\in\mathcal{N}:b\in\textnormal{{UAlg}}_{M,I}\text{ and }b\notin\textnormal{{Alg}}_{M,I}\},$
and $\textnormal{{Alg}}_{M,I}$ is the total value obtained by our (constrained) algorithm. Using the unconstrained benchmark, we have the basic identity
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;=\;\mathbb{E}_{M,I}\left[\textnormal{{UAlg}}_{M,I}\right]\;-\;\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right].$
We now bound the loss due to rejections. Partition buyers into those with moderate value and those with very large value:
$\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\right]$ $=\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}<2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$\qquad+\mathbb{E}_{M,I}\left[\sum_{b\in\textsc{Blocked}_{M,I}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$\leq 2\textnormal{{FracOpt}}_{\gamma}\cdot\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]$
$\qquad+\mathbb{E}_{M,I}\left[\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
Bounding the number of rejected buyers.
We begin by analyzing the first term, bounding the expected number of blocked buyers.
@@ 5 General Single-Minded Setting
We have
$\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]\;\leq\;\frac{1}{20}.$
Proof.
A buyer is rejected only if some item in their bundle is oversold. Thus
$|\textsc{Blocked}_{M,I}|\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}\mathbbm{1}{\left\{L_{M,I}(e)=k\right\}}\,(k-c(e)).$
Taking expectations and using linearity,
$\mathbb{E}_{M,I}\left[|\textsc{Blocked}_{M,I}|\right]\;\leq\;\sum_{e\in\mathcal{M}}\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\mathbb{P}\!\left[L_{M,I}(e)=k\right].$
By Lemma˜3.4, the random load $L_{M,I}(e)$ is upper bounded by a sum of negatively associated Bernoulli variables with total mean $\mu_{e}\leq c(e)/\gamma$. A standard Chernoff bound for negatively associated variables gives
Summing over $j$ gives
$\sum_{k=c(e)+1}^{\infty}(k-c(e))\,\Pr[L_{M,I}(e)=k]\;\leq\;\frac{1}{20m}.$
Summing over all $m$ items yields the bound. ∎
Plugging Lemma˜5.1 into our decomposition and using Lemma˜3.3, we obtain:
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{M,I}\right]\;\geq\;\frac{1}{40}\,\textnormal{{FracOpt}}_{\gamma}-\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right].$
The second term corresponds to buyers with extremely large valuations. We now control this contribution by mixing with a second, high-price mechanism.
Handling very large valuations.
Define an auxiliary mechanism $\textnormal{{Alg}}_{2}$ that posts a single menu entry consisting of all items, priced at $2\textnormal{{FracOpt}}_{\gamma}$. We first show that we may assume that extremely large valuations are rare.
Lemma 5.2.
$\mathbb{P}\!\left[\exists b:v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right]\leq 1/2.$
Proof.
If instead $\mathbb{P}\!\left[\exists b:v_{I}(b)\geq 2\textnormal{{FracOpt}}_{\gamma}\right]>1/2$, then simply posting one copy of the whole item set priced at $2\textnormal{{FracOpt}}_{\gamma}$ would yield expected value $\textnormal{{FracOpt}}_{\gamma}$, giving an $O(\gamma)$-approximation directly, which is already $O(m^{1/(B+1)})$. Thus we may assume the stated bound. ∎
We now show that $\textnormal{{Alg}}_{2}$ captures a constant fraction of the total contribution of very large valuations.
Lemma 5.3.
$\mathbb{E}_{M,I}\left[\textnormal{{Alg}}_{2}\right]\;\geq\;\frac{1}{2}\,\sum_{b\in\mathcal{N}}\mathbb{E}_{M,I}\left[v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right].$
Proof.
For fixed $M,I$,
[…]
Whenever all other buyers $b^{\prime}\neq b$ have $v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}$, the copy is available for $b$. Thus
$\textnormal{{Alg}}_{2}\;\geq\;\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\,\mathbbm{1}{\left\{\forall b^{\prime}\neq b:v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}\right\}}.$
Taking expectations and using independence of the $v_{b}$,
$\mathbb{E}_{I}\left[\textnormal{{Alg}}_{2}\right]$ $\;\geq\;\sum_{b}\mathbb{E}_{I}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\,\mathbb{P}\!\left[\forall b^{\prime}:v(b^{\prime})<2\textnormal{{FracOpt}}_{\gamma}\right]$
where the last inequality applies Lemma˜5.2. ∎
Combined algorithm and guarantee.
Let $\textnormal{{Alg}}_{1}$ denote the mechanism generated by the LP-based menu, and let the final algorithm Alg run $\textnormal{{Alg}}_{1}$ with probability $1/3$ and $\textnormal{{Alg}}_{2}$ with probability $2/3$. Combining the previous bounds, we obtain
$\mathbb{E}\left[\textnormal{{Alg}}\right]$ $\geq\frac{1}{3}\left(\frac{1}{40}\textnormal{{FracOpt}}_{\gamma}-\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\right)$
$\qquad+\frac{2}{3}\cdot\frac{1}{2}\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$=\frac{1}{120}\textnormal{{FracOpt}}_{\gamma}\;\geq\;\frac{\textnormal{{FracOpt}}}{120e(20m)^{1/(B+1)}}$
6 Graph Setting
In this section, we extend our framework to the setting where items correspond to edges and buyers correspond to routing requests in a general graph. Let $G=(V,E)$ be an undirected graph equipped with an integer capacity function $c:E\to\mathbb{Z}_{>0}$. Each buyer $b\in[n]$ is associated with a publicly known source-destination pair $(s_{b},t_{b})$. Upon arrival, buyer $b$ draws a value $v_{b}$ from a known distribution $\mathcal{D}_{b}$. Buyer $b$ derives value $v_{b}$ if and only if they receive a bundle of edges that contains a path connecting $s_{b}$ to $t_{b}$; otherwise their utility is zero. We proceed to prove Theorem˜1.3.
Let $\mathcal{P}_{s,t}$ be the set of all simple $s$-$t$ paths in $G$ and let $\mathcal{P}_{b}=\mathcal{P}_{s_{b},t_{b}}$. For each $b$, let $q_{b,v}=\mathbb{P}\!\left[v_{b}=v\right]$. We write $p\ni e$ to denote that a path $p$ uses edge $e$. As in the single-minded setting, we compare against the optimal prophet via a fractional relaxation. The LP now contains one variable for every possible routing choice:
$\max$ $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\cdot v$
s.t. $\sum_{b\in[n]}\sum_{v\in\mathcal{S}}\sum_{\begin{subarray}{c}p\in\mathcal{P}_{b}:\\ e\in p\end{subarray}}x_{b,v,p}\;\leq\;c(e),$ $\forall e\in E,$
$\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\;\leq\;q_{b,v},$ $\forall b\in[n],\,v\in\mathcal{S},$
$x_{b,v,p}\geq 0,$ $\forall b\in[n],\,v\in\mathcal{S},\,p\in\mathcal{P}_{b}.$
Let FracOpt denote the value of the optimal fractional LP solution.
Buyer Types and Generalized Bundles.
A key observation is that any $s$-$t$ path satisfies all buyers with the same source-target pair $(s,t)$. This motivates treating each pair $(s,t)$ as a type whose feasible bundles are exactly the $s$-$t$ paths $\mathcal{P}_{s,t}$.
We introduce a generalized bundle $B_{s,t}$, representing the act of serving an $(s,t)$-buyer. Each copy of $B_{s,t}$ will later be instantiated as a specific simple path $p\in\mathcal{P}_{s,t}$; different copies may correspond to different paths. The path associated with each copy is sampled before any buyer arrives, from a distribution over $\mathcal{P}_{s,t}$ that we define later.
For the moment, assume that every buyer $b$ with $(s_{b},t_{b})=(s,t)$ can only purchase copies of $B_{s,t}$ [^1] In Section D.2 we remove this assumption by proving that the menu prices are subadditive, so buyers never have an incentive to purchase a collection of tickets that indirectly covers their path instead of $B_{s,t}$ itself.. Under this assumption, buyers of the same type behave exactly like single-minded buyers for a generalized bundle $B_{s,t}$. We may therefore group them as a single demand class and randomize over how many copies of $B_{s,t}$ to create, just as we randomize over bundle copies in the single-minded setting.
Crucially, all arguments that relate the LP to the expected value of the unconstrained algorithm (Property 1) depend only on the total fractional mass assigned to each “bundle” and not on its internal structure. After grouping buyers by type in this way, the analysis of UAlg from the single-minded case carries over without change. What remains is to specify how we (i) construct the menu for each type $(s,t)$ and (ii) route buyers who purchase $B_{s,t}$ so that the load bounds (Property 2) are preserved at the edge level.
Menu Construction and Routing.
We now describe the construction of the menu $M$ in the graph setting. For each pair $(s,t)$ with at least one buyer, let
$\mathcal{A}_{s,t}\coloneqq\{b\in[n]:(s_{b},t_{b})=(s,t)\}$
be the set of buyers of type $(s,t)$, and define
$x_{(s,t),v}\;=\;\sum_{b\in\mathcal{A}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}}x_{b,v,p}\quad\text{and}\quad x_{(s,t),v,p}\;=\;\sum_{b\in\mathcal{A}_{s,t}}x_{b,v,p}.$
As in Section˜2.2, we can view $X^{(s,t)}=[x_{(s,t),v}]_{v\in\mathcal{S}}$ as the aggregate LP allocation over values for type $(s,t)$. Let $w_{s,t}$ be the important value of this vector, defined exactly as in the single-minded case. Depending on whether $w_{s,t}$ is tight or crucial, we apply the same three-case rule as in (Step 2) to determine how many copies of $B_{s,t}$ to post at price $w_{s,t}$ and at price $w_{s,t}+1$.
Formally, the menu for type $(s,t)$ consists of $c_{s,t}$ copies of $B_{s,t}$, each priced either at $w_{s,t}$ or $w_{s,t}+1$, according to the usual case distinction from Section˜3.1. This fully specifies the prices and multiplicities of the generalized bundles.
We now define how each copy of $B_{s,t}$ is instantiated as a concrete $s$-$t$ path. For every copy of $B_{s,t}$ priced at $w$, we independently sample a path $p\in\mathcal{P}_{s,t}$ using roulette drawing with probabilities proportional to the LP mass at value $w$ (we omit the $(s,t)$ index, since we fix an $(s,t)$ pair):
$p\sim\frac{x_{w,p}}{\sum_{p^{\prime}\in\mathcal{P}_{s,t}}x_{w,p^{\prime}}}.$
Similarly, for every copy of $B_{s,t}$ priced at $w+1$, we independently sample
$p\sim\frac{\sum_{v>w}x_{v,\,p}}{\sum_{p^{\prime}\in\mathcal{P}_{s,t}}\sum_{v>w}x_{v,\,p^{\prime}}}.$
This completes the definition of the menu and routing policy in the graph setting.
Load Analysis.
We now show that the expected load induced by this construction satisfies the same kind of bound as in Lemma˜3.4, yielding Property 2 for graphs.
For a fixed type $(s,t)$ and edge $e$, let $L_{s,t}(e)$ denote the total load on edge $e$ due to buyers of type $(s,t)$ under the (unconstrained) execution of the menu. Let $P_{s,t}^{e}\subseteq\mathcal{P}_{s,t}$ be the set of $s$-$t$ paths that use edge $e$. The following lemma is proved using a similar logic as in the single-minded case. A formal proof of the lemma is deferred to the Section˜D.1.
Lemma 6.1 (Property 2).
For every edge $e\in E$ and any scaling factor $\gamma>1$, there exist $\{0,1\}$ negatively associated random variables $Z_{1},\ldots,Z_{k}$ such that
$L_{M,I}(e)\leq\sum_{i=1}^{k}Z_{i}=Z_{M,I}(e)\quad\text{and}\quad\mathbb{E}_{M,I}\left[L_{M,I}(e)\right]\leq\mathbb{E}_{M,I}\left[Z_{M,I}(e)\right]\leq\frac{c(e)}{\gamma}.$
Combining Lemma˜6.1 with the value guarantee for the unconstrained algorithm (Lemma˜3.3), and with the same blocking analysis used in the single-minded setting, we obtain the claimed $O\bigl(m^{1/(B+1)}\bigr)$ competitive ratio for the graph-routing model.
7 Lower Bounds
We now formally prove our lower bounds for the general capacity setting, i.e. Theorem˜1.4 and Theorem˜1.5. The refined lower bounds for the unit capacity setting, i.e. Theorem˜1.6 and Theorem˜1.7 are deferred to Appendix˜E.
7.1 Proof of Theorem 1.4
It suffices to prove the theorem under the assumption $1\leq B<\ln m$, since for $B\geq\ln m$ the trivial $\Omega(1)$ lower bound already yields the claim. We recall the notion of $r$-way qualitatively independent (QI) partitions due to Rényi [renyi].
Definition 7.1 ($r$-way qualitatively independent partitions).
Let $U$ be an $n$-element ground set. A partition of $U$ into $t$ classes is a collection $\mathcal{P}=\{P_{1},\dots,P_{t}\}$ of disjoint sets whose union is $U$. A family of partitions
$\mathcal{F}=\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(k)}\},\qquad\mathcal{P}^{(j)}=\{P^{(j)}_{1},\dots,P^{(j)}_{t}\}$,
is called $r$-way qualitatively independent if for every choice of $r$ distinct indices $j_{1},\dots,j_{r}\in[k]$ and every choice of classes $a_{1},\dots,a_{r}\in[t]$ we have
$P^{(j_{1})}_{a_{1}}\cap\dots\cap P^{(j_{r})}_{a_{r}}\neq\emptyset$.
Let $g(n,t,r)$ denote the largest integer $k$ for which such a family of $k$ partitions of an $n$-element set into $t$ classes exists. Poljak and Tuza [poljak] proved the following lower bound.
Theorem 7.2 (Theorem 5 in [poljak]).
For all integers $n,t,r\geq 1$,
$g(n,t,r)\;\geq\;\frac{er}{t}\cdot\exp\!\left(\frac{n}{rt^{r}}\right).$
We will apply this bound with $n$ equal to the number of items $m$ and $r=B+1$. The key step is to choose the number of classes $t$ so that $g(m,t,r)$ is approximately $Bt^{t}$. In Appendix G we prove the following.
Lemma 7.3.
For every integer $m\geq 16$ and every integer $B$ with $1\leq B<\ln m$, define

### v2

@@ 6.1 The Unconstrained Algorithm and Its Properties
Whenever all other buyers $b^{\prime}\neq b$ have $v_{b^{\prime}}<2\textnormal{{FracOpt}}_{\gamma}$, the posted copy of $\mathcal{M}$ is available for $b$. Thus,
$\textnormal{{Alg}}_{2}\;\geq\;\sum_{b\in\mathcal{N}}v_{b}\,\mathbbm{1}{\left\{v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\,\mathbbm{1}{\left\{\forall b^{\prime}\neq b:v_{b^{\prime}}<2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\,.$
Taking expectations and using independence of the realized values of different buyers,
$\mathbb{E}_{I}\left[\textnormal{{Alg}}_{2}\right]$ $\;\geq\;\sum_{b}\mathbb{E}_{I}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\,\mathbb{P}\!\left[\forall b^{\prime}\neq b:v_{b^{\prime}}<2\,\textnormal{{FracOpt}}_{\gamma}\right]$
where the last inequality holds due to the definition of Case 2. ∎
Combined Algorithm and Competitive Ratio.
Let $\textnormal{{Alg}}_{1}$ denote the mechanism generated by the EA-LP-based menu, and let the final algorithm Alg run $\textnormal{{Alg}}_{1}$ with probability $1/3$ and $\textnormal{{Alg}}_{2}$ with probability $2/3$. Combining the previous bounds, we conclude the proof of Theorem 2.2:
$\mathbb{E}\left[\textnormal{{Alg}}\right]$ $\geq\frac{1}{3}\left(\frac{\textnormal{{FracOpt}}_{\gamma}}{40}-\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\right]\right)+\frac{2}{3}\cdot\frac{1}{2}\sum_{b}\mathbb{E}\left[v_{b}\mathbbm{1}{\left\{v_{b}\geq 2\,\textnormal{{FracOpt}}_{\gamma}\right\}}\right]$
$=\frac{\textnormal{{FracOpt}}_{\gamma}}{120}\;\geq\;\frac{\textnormal{{FracOpt}}}{120e(20m)^{1/(B+1)}}.$
9 Online Network Routing: The Proof of Theorem 2.3
We next extend our framework to the slightly more general setting where items correspond to edges and buyers correspond to routing requests in a network. Let $G=(V,E)$ be directed network equipped with an integer capacity function $c:E\to\mathbb{Z}_{>0}$. Each buyer $b\in[n]$ is associated with a publicly known source-target pair $(s_{b},t_{b})$. Upon arrival, buyer $b$ draws a value $v_{b}$ from a known distribution $\mathcal{D}_{b}$. Buyer $b$ derives value $v_{b}$ if they receive bundle of edges that contains a path connecting $s_{b}$ to $t_{b}$; otherwise their value is zero.
Let $\mathcal{P}_{s,t}$ be the set of all simple $s$-$t$ paths in $G$ and let $\mathcal{P}_{b}=\mathcal{P}_{s_{b},t_{b}}$. As above, for each $b$, we let $q_{b,v}=\mathbb{P}\!\left[v_{b}=v\right]$. We write $p\ni e$ to denote that a path $p$ uses edge $e$.
As in the single-minded setting, we compare against the optimal prophet via the following fractional relaxation. The EA-LP now contains one variable for every possible routing choice:
$\max$ $\sum_{b\in[n]}\sum_{v\in\mathcal{U}}\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\cdot v$
s.t. $\sum_{b\in[n]}\sum_{v\in\mathcal{U}}\sum_{p\in\mathcal{P}_{b}:e\in p}x_{b,v,p}\;\leq\;c(e),$ $\forall e\in E,$
$\sum_{p\in\mathcal{P}_{b}}x_{b,v,p}\;\leq\;q_{b,v},$ $\forall b\in[n],\,v\in\mathcal{U},$
$x_{b,v,p}\geq 0,$ $\forall b\in[n],\,v\in\mathcal{U},\,p\in\mathcal{P}_{b}.$
As before, we solve (and use) a $\gamma$-scaled down of EA-LP, with $c(e)/\gamma$, instead of $c(e)$, on the rhs of the first constraint. We let FracOpt (resp. $\textnormal{{FracOpt}}_{\gamma}$) denote the value of the optimal fractional (resp. scaled down) EA-LP solution.
Buyer Types and Generalized Bundles.
A key observation is that any $s$-$t$ path satisfies all buyers with the same source-target pair $(s,t)$. This motivates treating each pair $(s,t)$ as a type whose feasible bundles are all $s$-$t$ paths in $\mathcal{P}_{s,t}$. We introduce the notion of a generalized bundle $B_{s,t}$, representing the act of serving an $(s,t)$-buyer. Each copy of $B_{s,t}$ is instantiated as a specific path $p\in\mathcal{P}_{s,t}$; different copies of bundle $B_{s,t}$ may correspond to different sts-t paths. The path associated with each copy is sampled before any buyer arrives from a distribution over $\mathcal{P}_{s,t}$ defined below.
For simplicity, we assume that every buyer $b$ with $(s_{b},t_{b})=(s,t)$ can only purchase copies of $B_{s,t}$ [^1] In Appendix D.2, we remove this assumption by proving that the menu prices are subadditive; so the buyers never have an incentive to purchase a collection of paths whose union includes an sts-t path, instead of purchasing a copy of $B_{s,t}$.. Then, buyers of the same type behave like single-minded buyers w.r.t. a generalized “bundle” $B_{s,t}$. We may therefore group them as a single demand type and randomize over how many copies of $B_{s,t}$ to create, just as we randomize over bundle copies in the single-minded setting.
Crucially, all arguments that relate FracOpt to the expected value of the unconstrained algorithm (Property 1, Lemma 6.1) depend only on the total fractional allocation to each generalized “bundle” $B_{s,t}$ and not on the bundle’s internal structure (i.e., how different copies of $B_{s,t}$ are routed through different sts-t paths). After grouping buyers by type, the analysis of UAlg from the single-minded case carries over without any change. It remains to specify how (i) we construct the menu for each $(s,t)$ type; and (ii) we route buyers who purchase $B_{s,t}$ so that the load bounds (Property 2, Lemma 6.3) are preserved at the edge level.
Menu Construction and Routing.
@@
Formally, the menu for type $(s,t)$ consists of $c_{s,t}$ copies of $B_{s,t}$, each priced either at $w_{s,t}$ or $w_{s,t}+1$, according to the case distinction in Algorithm 1. This fully specifies the number of copies offered at each price for each generalized bundle $B_{s,t}$.
We now define how each copy of $B_{s,t}$ is instantiated as a concrete $s$-$t$ path. For every copy of $B_{s,t}$, we independently sample a path $p\in\mathcal{P}_{s,t}$ using roulette drawing with probability proportional to EA-LP allocation (we omit the $(s,t)$ index, because the $(s,t)$ pair is fixed):
$\mathbb{P}\!\left[\mbox{path $p$ is chosen}\right]=\frac{\sum_{v\geq w}x_{v,\,p}}{\sum_{p^{\prime}\in\mathcal{P}_{s,t}}\sum_{v\geq w}x_{v,\,p^{\prime}}}.$
This completes the definition of the menu and the routing in the online network routing setting.
Edge Load.
We now show that the expected load induced by this construction satisfies the same kind of bound as in Lemma 6.3, yielding Property 2 for the load of network edges.
For a fixed type $(s,t)$ and edge $e$, let $L_{s,t}(e)$ denote the total load on edge $e$ due to buyers of type $(s,t)$ under the (unconstrained) execution of online algorithm with the path allocation menu defined above. Let $P_{s,t}^{e}\subseteq\mathcal{P}_{s,t}$ be the set of $s$-$t$ paths that use edge $e$. The following lemma can proved similarly to Lemma 6.3 (the proof is deferred to Appendix D.1).
Lemma 9.1.
For every edge $e\in E$, the path allocation of the unconstrained algorithm with the routing menu guided by the $\gamma$-scaled down EA-LP and constructed as above satisfies:
$\mathbb{E}_{M,I}\left[L_{s,t}(e)\right]\;\leq\;\mathbb{E}_{M,I}\left[Z_{s,t}(e)\right]\;\leq\;\sum_{b\in\mathcal{N}_{s,t}}\sum_{v\in\mathcal{U}}\sum_{p\in P_{s,t}^{e}}x_{b,v,p}\;\leq\;\frac{c(e)}{\gamma},$
where $Z_{s,t}(e)$ is a sum of negatively associated random variables.
Combining Lemma 9.1 with the expected social welfare guarantee for the unconstrained algorithm of Lemma 6.1, and using the value of $\gamma$ and the analysis used for the general single-minded setting in Section 8, we obtain a competitive ratio of $O(m^{1/(B+1)})$ for online network routing (Theorem 2.3).
10 Lower Bounds
Next, we prove Theorem 2.4 and give a proofsketch of Theorem 2.5. The full proof of Theorem 2.5 along with refined lower bounds for the case of unit capacities are presented in Appendix E.
10.1 The Proof of Theorem 2.4
It suffices to consider the case where $1\leq B<\ln m$ (if $B\geq\ln m$, the trivial $\Omega(1)$ bound implies the claim). We use $r$-way qualitatively independent (QI) partitions introduced by Rényi (1970).
Definition 10.1 ($r$-way QI partitions).
Let $U$ be an $n$-element set. A partition of $U$ into $t$ classes is $\mathcal{P}=\{P_{1},\dots,P_{t}\}$, where the different classes are mutually disjoint and their union is $U$. A family $\mathcal{F}=\{\mathcal{P}^{(1)},\dots,\mathcal{P}^{(k)}\}$ of $t$-partitions is $r$-way qualitatively independent if for every choice of $r$ distinct partition indices $j_{1},\dots,j_{r}\in[k]$ and $r$ class indices $a_{1},\dots,a_{r}\in[t]$, one for each partition, it holds that $P^{(j_{1})}_{a_{1}}\cap\cdots\cap P^{(j_{r})}_{a_{r}}\neq\emptyset$. Let $g(n,t,r)$ be the maximum size of such a family.
Poljak and Tuza (1989) proved the following.
Theorem 10.2 ((Poljak and Tuza, 1989, Theorem 5)).
For all integers $n,t,r\geq 1$,
@@ 10 Lower Bounds
In the following, we set $n=m$ and $r=B+1$, and choose $t$ so that $g(m,t,B+1)\geq B\,t^{t}$. The proof is by direct substitution into Theorem 10.2 and is deferred to Appendix F.
@@
For every $m\geq 16$ and $1\leq B<\ln m$, define